# A New Census of the Universe’s Entropy

### v1

@@ Abstract
The question of what is the total entropy of the universe, how it compares to the maximal entropy of de Sitter space, and how it is distributed across the universe’s components, bears considerable importance for a number of reasons. Here, we first update the computation of the entropy associated with various sectors of the observed universe, including in the diffuse cosmic and late-time gamma-ray and neutrino backgrounds, in baryonic matter both in diffuse components, in stars and stellar remnants, and in cosmic rays; we then update, crucially, the estimate of entropy in stellar-mass and super-massive black holes, whose abundance and mass function has come into increasingly sharp definition with recent observations and with the rapidly growing statistics of black-hole-black-hole mergers observed with gravity wave detectors. We also provide a new, corrected estimate of the potential entropy associated with a stochastic gravitational wave background, with dark sector radiations, and with several dark matter models. Finally, we utilize the similarly recently updated constraints on the abundance of hypothetical primordial black holes – black holes, that is, of non-stellar origin – to assess the maximal amount entropy they could store. We find that if supermassive primordial black holes exist, they can dominate the entropy budget of the universe consistently with current constraints on their abundance and mass function, to a level potentially not distant from the posited entropy associated with the cosmic event horizon of de Sitter spacetime. The same conclusion holds for certain dark sector models featuring a large number of dark degrees of freedom.
@@ 1 Introduction
Crucially, the cosmic event horizon’s entropy, associated with the de Sitter horizon in a cosmological-constant-dominated universe, can provide [^1] Note that it is at present a matter of active discussion whether or not the de Sitter cosmological horizon entropy constitutes an “actual” entropy that acts, that is, like black holes’ horizon entropy, and if these arguments can be applied to horizons more broadly defined (see Ref. [18] for extension of some black hole horizon properties to de Sitter horizons, see Ref. [19] for analysis of thermodynamic status of de Sitter horizons, see Ref. [20] for extension to general causal horizons) – for instance, it is not necessarily common to consider Unruh horizons and resulting Unruh radiation as adding entropy differentially for observers accelerating at different rates, though this has been analyzed in Ref. [20] – and it is not agreed upon whether or not cosmological horizon entropy should be considered as part of the universe’s total entropy, an upper bound, or ignored entirely (for instance, Ref. [11] argues that when considering a comoving volume and assuming homogeneity, one can safely infer entropy densities beyond the horizon, and so the horizon entropy need not be considered). an upper bound on the universe’s entropy content [11, 21]. The entropy of the cosmic event horizon is, as it should be, much larger than the estimate of the combined contributions of internal components, marking it a pivotal element in understanding thermodynamic limits and cosmological models [22]. Studies suggest that the entropy associated with the cosmic event horizon dominates the entropy budget of the universe since early cosmological epochs [21].
@@
Understanding cosmic entropy is not a mere academic exercise; it bears profound implications for cosmological models. Accurately determining entropy provides a window into cosmic evolution, offering insights into the irreversible processes that shape the universe, the arrow of time, and constraints on models like $\Lambda$CDM. Evaluating the cosmological entropy budget enables insights into the nature of thus-far mysterious components of the universe’s energy density makeup, such as dark matter, dark energy, possible dark-sector relativistic species, diffuse gravitational backgrounds, and the origin of cosmic acceleration, as well as a deeper theoretical understanding of cosmology rooted in thermodynamics [13, 24].
@@ 2 Cosmic Event Horizon
The entropy associated with de Sitter (dS) and quasi-de Sitter (q-dS) spacetimes has been widely researched in a number of frameworks, including semiclassical thermodynamics, string theory, holography, and quantum gravity. The foundational work of Gibbons and Hawking established that de Sitter spacetime possesses an entropy $S{\text{dS}}=A/4G$, proportional to the area of the cosmological horizon, and associated with a temperature $T=H/2\pi$ defined by the Hubble parameter $H$ [21]. This entropy parallels what is well known in black hole thermodynamics, and is interpreted as a measure of quantum states accessible within the dS causal patch [25, 26, 27, 28, 29]. However, unlike black hole entropy, the microscopic origin of $S{\text{dS}}$ remains a relatively contentious topic, with various proposals ranging from quantum soft modes and finite-dimensional Hilbert spaces to string-theoretic mechanisms and holographic duals [30, 31, 32, 33, 34].
Holographic interpretations, particularly through the dS/CFT correspondence, have been a primary approach to understanding de Sitter entropy and its implications for quantum gravity. This conjecture, proposed by Strominger [25], suggests that quantum gravity in dS spacetime is dual to a Euclidean CFT on a spacelike boundary at $\mathcal{I}^{+}$. Bousso and collaborators have connected aspects of dS entropy to conformal field theory measures, including the Cardy formula, but the dual CFT often exhibits non-unitary properties, raising concerns about its physical completeness [35, 36, 37]. The difficulty in defining a boundary theory for dS, as opposed to AdS holography, has led to introduce modifications, including quantum-deformed symmetry groups, entropy islands, and extremal surfaces, particularly in lower-dimensional settings [38, 39, 40, 41]. While these approaches partially recover the Gibbons-Hawking entropy and provide hints about its microscopic structure, the gap between semiclassical and fully holographic descriptions remains a major challenge.
In string theory, several studies suggest that de Sitter entropy originates from microscopic degrees of freedom associated with strings or branes near the cosmological horizon. For instance, Halyo suggested that $S_{\text{dS}}$ can be understood as the entropy of a highly excited string at the Hagedorn temperature, renormalized due to the gravitational redshift near the horizon [32, 37]. Other approaches, such as those by Sanchez and Silverstein, connect string density of states or compactification flux vacua to the entropy-temperature relationship of de Sitter spacetime, finding consistency with semiclassical predictions but highlighting challenges with constructing stable dS vacua in string theory due to no-go theorems or metastability [30, 42, 43, 34]. These efforts have advanced understanding of the connection between quantum gravity and cosmological spacetimes, but the instability of dS in string theory often limits their applicability [31, 44].
Thermodynamic interpretations of dS and q-dS spacetimes also reveal deep connections between entropy and cosmology. In the context of slow-roll inflation, deviations from exact de Sitter symmetry influence the horizon area and entropy variations due to scalar field fluctuations and quantum diffusion [26, 45, 46]. Entropy bounds appear to impose constraints on inflationary e-foldings and observable modes, reflecting the finite information capacity of de Sitter horizons during inflationary epochs [26, 46]. Stochastic approaches to inflation further link entropy changes to coarse-grained field dynamics and long-wavelength perturbations, demonstrating connections between IR effects and the overall thermodynamic behavior of quasi-de Sitter horizons [45]. These insights not only refine thermodynamic predictions for inflationary cosmology but also suggest entropy as a fundamental measure of spacetime structure.
Further complexities arise in spacetime configurations where black hole horizons coexist with de Sitter horizons, such as Schwarzschild-de Sitter and Nariai spacetime geometries. In such cases, the entropy contributions from the black hole and cosmological horizons evolve dynamically, raising questions about global thermodynamic consistency [27, 47, 48]. String-based models have been used to study thermodynamic and holographic connections in these setups, drawing parallels to entropy competition in AdS black holes and extending metrics to include horizon interactions through localized degrees of freedom [32, 49, 48]. These scenarios underscore the intricate relationship between horizon entropy, thermodynamics, and causal structure in de Sitter spacetimes.
The lack of a unified microscopic theory for de Sitter entropy limits understanding of quantum-gravitational phenomena in dS and q-dS universes [30, 50, 34]. Proposals such as species bounds, quantum extremal surfaces, and q-deformed holographic duals provide promising tools but require further development [38, 39, 50]. Moreover, the interplay between entropy corrections, such as logarithmic terms derived from quantum effects, and the semiclassical entropy formula suggests that deviations at high curvatures or in quantum regimes could provide new insights [44, 43, 28, 29].
@@
Here, we resort to the original Gibbons-Hawking form $S{\text{dS}}=A/4G$ and update the entropy associated with the cosmic event horizon (CEH) with the latest central values of cosmological parameters given in Ref. [51]. Using the approach of Ref. [11], and marginalizing over the uncertainty in cosmological parameters, we find
[…]
@@ 3 Photon and neutrino diffuse backgrounds
$s_{\rm CMB}=\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{c^{3}\hbar^{3}}g_{\gamma}T_{\rm CMB}^{3}=1.476\times 10^{9}~{}k_{B}~{}\mathrm{m^{-3}},$ (3.1)
with $g_{\gamma}=2$ above, and a completely negligible error from the error on the CMB temperature. The corresponding CMB entropy in the universe reads
$S_{\rm CMB}=s_{\rm CMB}V_{\rm obs}=(5.196\pm 0.16)\times 10^{89}\ k_{B},$ (3.2)
@@
$s_{\rm EBL}$ $\simeq$ $8.6\times 10^{6}~{}k_{B}~{}\mathrm{m^{-3}},\qquad S_{\rm EBL}=s_{\rm EBL}V_{\rm obs}=3.02\times 10^{87}\ k_{B},$ (3.3)
@@
$s_{{\rm C}\nu{\rm B}}=\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{c^{3}\hbar^{3}}g_{\nu}T_{{\rm C}\nu{\rm B}}^{3}=1.408\times 10^{9}~{}k_{B}~{}\mathrm{m^{-3}}.$ (3.5)
with $g_{\nu}=6$ in the equation above for the 6 internal degrees of freedom of the three Standard Model neutrinos, and the (7/8) coefficient is the standard Fermi-Dirac correction. The error corresponding to the cosmic neutrino temperature is, as for the CMB, completely negligible (one in a billion). The corresponding entropy in the universe reads
@@
$\langle v\rangle\simeq 160\ (1+z)\ \frac{\rm km}{\rm sec}\ \left(\frac{\rm eV}{m_{\nu}}\right).$ (3.7)
@@
$\lambda_{\rm fs}\simeq 2\ {\rm Mpc}\left(\frac{\rm eV}{m_{\nu}}\right).$ (3.9)
@@ 4.1 Diffuse unbound baryons
The WHIM is the primary reservoir of baryonic matter ($\sim$40-50of baryons, strongly supported by both simulations [82, 83, 84] and observations [85, 86, 87] and resides in filamentary structures of the cosmic web, with temperatures of 105-107 K [82, 83, 84, 88, 89]. UV/X-ray absorption studies (e.g., OVI, OVII, OVIII) confirm WHIM’s presence [85, 86, 90]. Stacking techniques ([87]) and thermal Sunyaev-Zeldovich (SZ) surveys ([91]) reveal WHIM in dense filaments, though detection of hotter phases ($>10^{6}$ K) remains incomplete.
The diffuse IGM contains $\sim$30-40of baryons at lower temperatures, while galaxies and ISM ($\sim$10%) and ICM ($\sim$4-5%) are much smaller contributors [85, 83, 92, 93]. This distribution aligns with $\Lambda$CDM predictions and resolves much of the “missing baryon” problem. We note that distinctions between IGM and WHIM (near $10^{5}$ K threshold) remain blurred in some studies [85, 92]. Additionally, detecting the full baryon fraction in WHIM across all density environments is limited by current observational sensitivities [86, 87].
@@
ISM (Cold ISM) $10^{1}-10^{2}$ ($10-100$) $\sim 65$ $(5.7-57)\times 10^{80}$
ISM (Warm ISM) $\sim 10^{4}$ ($0.1-1$) 75-76 $(6.6-67)\times 10^{80}$
@@
We estimate the contribution to the total entropy of the universe of diffuse baryons using the ranges of density and temperature detailed upon above; Tab. 1 lists the entropy per baryon and the total entropy for each component, indicating that all components have comparable entropy contributions, with the IGM components contributing most likely the largest part to the entropy in diffuse backgrounds, roughly a factor 3 more than the ISM and 30 or more than the ICM. Overall, we find that, at most, diffuse baryons contribute to the entropy of the universe an amount
@@ 4.2 Baryons in stellar and compact objects
[…]
@@
The stellar remnants formed from these stars are predominantly white dwarfs, which are the endpoints of stars with initial masses less than $\sim 8$ $10M_{\odot}$ [107, 102]. The Milky Way is estimated to contain approximately $\sim 10^{9}-10^{10}$ white dwarfs, distributed across a range of cooling stages. Recent cooling models, accounting for core processes such as latent heat from crystallization and energy release during ${}^{22}\mathrm{Ne}$ sedimentation, suggest that older, “cold” white dwarfs ($\sim 1\times 10^{5}\,\mathrm{K}$) dominate the population, likely comprising $>90\%$ of all white dwarfs [102, 108, 109]. This implies a cold white dwarf population on the order of $2-3\times 10^{21}$ in the universe, while younger, hot white dwarfs ($\sim 5\times 10^{8}\,\mathrm{K}$) constitute a smaller fraction (1-3%), totaling $\sim 10^{20}$ objects [102, 108, 109].
Neutron stars, formed via supernova explosions of stars in the $8-25M_{\odot}$ range, are significantly less common than white dwarfs due to the IMF bias toward lower-mass progenitors [99, 104]. Population synthesis models predict approximately $\sim 10^{8}$ neutron stars in the Milky Way, primarily cooling remnants at temperatures of $\sim 10^{7}\,\mathrm{K}$ [110, 111]. Scaling to the universe, the cold neutron star population is on the order of $\sim 2-3\times 10^{20}$. Conversely, newly formed, “hot” neutron stars $(\sim 10^{9}\,\mathrm{K})$ exist only briefly (for seconds to minutes post-supernova) and number $<10^{3}$ in the Milky Way at any given time, extrapolating to $\sim 10^{15}$ in the observable universe [110, 112].
@@
MS ($\sim 1.5M_{\odot}$) $\sim$1-5 billion $\sim 3-10\times 10^{20}$ 3,718 (1.8-5.9)×$\times 10^{81}$
MS ($\sim 7M_{\odot}$) $\sim$10 million $\sim 2-5\times 10^{17}$ 4,671 (1.5 -3.7)×$\times 10^{78}$
MS ($\sim 11M_{\odot}$) $\sim$1 million $\sim 2-5\times 10^{16}$ 4,331 (1.4 - 3.5)×$\times 10^{77}$
MS ($\sim 25M_{\odot}$) $\sim$10,000 $\sim 2-5\times 10^{13}$ 5,173 (1.7 -4.1 )×$\times 10^{74}$
Hot WD $\sim$50-100 million $\sim 1-3\times 10^{20}$ 5.72 (0.9 -2.7)×$\times 10^{78}$
Cold WD $\sim$8-9 billion $\sim 2-3\times 10^{21}$ 1.86×$\times 10^{-5}$ (6.0 -8.9)×$\times 10^{73}$
Hot NS $\sim$10-1,000 $\sim 2-5\times 10^{15}$ 5.45×$\times 10^{-3}$ (1.7 -4.3)×$\times 10^{70}$
Cold NS $\sim$100 million $\sim 2-3\times 10^{20}$ 5.45×$\times 10^{-5}$ (1.7 -2.6)×$\times 10^{73}$

### v2

@@ Abstract
The question of what is the total entropy of the universe, how it compares to the maximal entropy of de Sitter space, and how it is distributed across the universe’s components, bears considerable importance for a number of reasons. Here, we first update the computation of the entropy associated with various sectors of the observed universe, including in the diffuse cosmic and late-time gamma-ray and neutrino backgrounds, in baryonic matter both in diffuse components, in stars and stellar remnants, and in cosmic rays; we then update, crucially, the estimate of entropy in stellar-mass and super-massive black holes, whose abundance and mass function has come into increasingly sharp definition with recent observations and with the rapidly growing statistics of black-hole-black-hole mergers observed with gravitational wave detectors. We also provide a new, corrected estimate of the potential entropy associated with a stochastic gravitational wave background, with dark sector radiations, and with several dark matter models. Finally, we utilize the similarly recently updated constraints on the abundance of hypothetical primordial black holes – black holes, that is, of non-stellar origin – to assess the maximal amount of entropy they could store. We find that if supermassive primordial black holes exist, they can dominate the entropy budget of the universe consistently with current constraints on their abundance and mass function, to a level potentially not distant from the posited entropy associated with the cosmic event horizon of de Sitter spacetime. The same conclusion holds for certain dark sector models featuring a large number of dark degrees of freedom.
@@ 1 Introduction
Crucially, the cosmic event horizon’s entropy, associated with the de Sitter horizon in a cosmological-constant-dominated universe, can provide an upper bound on the universe’s entropy content [11, 18]. The entropy of the cosmic event horizon is, as it should be, much larger than the estimate of the combined contributions of internal components, marking it a pivotal element in understanding thermodynamic limits and cosmological models [19]. Studies suggest that the entropy associated with the cosmic event horizon dominates the entropy budget of the universe since early cosmological epochs [18].
@@
[…]
@@
A note on uncertainties: in our estimates below, whenever we quote range, actual confidence-level ranges are not available. It is in many cases discussed in what follows extremely problematic to quote confidently, or to distinguish and estimate, systematic and statistical uncertainties. The ranges are rather derived and quoted from careful literature surveys. There are exceptions, such as the calculation of the entropy associated with the CMB or with the cosmic neutrino background. In those cases, we quote uncertainties with the $\pm$ notation. In some cases, for brevity, we indicate ranges in order of magnitude in the exponent of powers of 10. In all cases, we indicate in detail the references used to extract the quoted rates.
Throughout this work, we use SI units, or occasionally astronomical units, in all equations and results to ensure clarity and facilitate comparison across disciplines; note that however, sometimes, for simplicity we revert to natural units ($c=\hbar=k_{B}=1$).
@@ 2 Cosmic Event Horizon
The entropy associated with de Sitter (dS) and quasi-de Sitter (q-dS) spacetimes has been established through the foundational work of Gibbons and Hawking, who showed that de Sitter spacetime possesses an entropy proportional to the area of the cosmological horizon, associated with a temperature proportional to the Hubble parameter [18]. This entropy parallels what is well known in black hole thermodynamics, and is interpreted as a measure of quantum states accessible within the dS causal patch [22, 23, 24, 25].
For the purposes of this paper, we consider a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) cosmological model with a positive cosmological constant. In this context, the CEH is defined as the boundary of the region from which light signals can reach a given observer in the infinite future. The proper distance from us to the present location of the cosmological horizon on a constant time slice is denoted by $R$, with the horizon area $A=4\pi R^{2}$ and associated Bekenstein-Hawking entropy $S_{\text{CEH}}=A/(4G)$.
The generalized second law of thermodynamics (GSL) for cosmological horizons states that the Bekenstein-Hawking entropy of the CEH plus the coarse-grained entropy inside of it does not decrease. This GSL has been proved under robust assumptions for perturbations of empty de Sitter space [26]. While an evolving universe represents more than an infinitesimal perturbation of de Sitter spacetime, the background growth of horizon area means that only an exceptionally large number of species could boost the matter content sufficiently to make matter entropy compete with horizon entropy growth.
In the following sections, we calculate the entropy contributions from various components of the universe and assess whether the total matter entropy could approach levels that would challenge the GSL under extreme assumptions about dark matter properties.
@@
Here, we resort to the original Gibbons-Hawking form $S_{\text{dS}}=\frac{k_{B}c^{3}A}{4G\hbar}$ and update the entropy associated with the CEH with the latest central values of cosmological parameters given in Ref. [27]. Using the approach of Ref. [11], and marginalizing over the uncertainty in cosmological parameters, we find
$R_{\rm CEH}=a(t_{now})\int_{t=t_{now}}^{\infty}\frac{c}{a(t)}\,dt=(1.577\pm 0.009)\times 10^{26}\ {\rm m}=16.68\pm 0.09~\mathrm{Glyr}.$ (2.1)
@@
Note that throughout this work, we assume that the present accelerated expansion of the Universe is driven by a true cosmological constant, corresponding to a constant vacuum energy density with equation of state w=1w=-1, which remains unchanged into the far future. This assumption implies that the Universe asymptotically evolves toward a de Sitter phase, and consequently, that a finite cosmological event horizon forms, characterized by radius $R_{\mathrm{CEH}}=H_{\Lambda}^{-1}$ and an associated Gibbons-Hawking entropy. If, instead, dark energy were to decay or transition into another dynamical phase, the CEH might not exist or could become infinitely large, thus modifying or eliminating its entropy contribution. The finite CEH entropy computed in this work therefore relies crucially on the assumption of a persistent cosmological constant.
@@ 3 Photon and neutrino diffuse backgrounds
$s_{\rm CMB}=\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{c^{3}\hbar^{3}}g_{\gamma}T_{\rm CMB}^{3}={\color[rgb]{0,0,0}(1.476\pm 0.001)}\times 10^{9}~k_{B}~\mathrm{m^{-3}},$ (3.1)
with $g_{\gamma}=2$ above. The corresponding CMB entropy in the universe reads
$S_{\rm CMB}=s_{\rm CMB}V_{\rm obs}=({\color[rgb]{0,0,0}5.20}\pm 0.16)\times 10^{89}\ k_{B},$ (3.2)
@@
$s_{\rm EBL}$ $\simeq$ ${\color[rgb]{0,0,0}(6.6\pm 0.2)}\times 10^{6}~k_{B}~\mathrm{m^{-3}},\qquad S_{\rm EBL}=s_{\rm EBL}V_{\rm obs}={\color[rgb]{0,0,0}(2.3\pm 0.1)}\times 10^{87}\ k_{B},$ (3.3)
@@
In greater detail, we find that by far the largest contributor to the EBL entropy is the infrared emission, primarily produced by starlight re-scattered by dust, which entirely encompasses the EBL entropy presented above; the optical (OP) and UV-and-higher-frequency (HE) contributions are:
$S_{\rm OP}\simeq{\color[rgb]{0,0,0}(2.70\pm 0.06)}\times 10^{85}\ k_{B},\quad S_{\rm HE}\simeq{\color[rgb]{0,0,0}(6.1\pm 0.1)}\times 10^{79}\ k_{B}.$ (3.4)
Interestingly, it has been claimed by Bousso and collaborators in Ref. [39] that it is important to determine the most significant sources of late-time matter entropy production (excluding horizon entropies), because high entropy production can be used as a proxy for observers when analyzing anthropic selection effects on $\rho_{\Lambda}$, the cosmological constant. Their argument is that while the landscape of possible values for $\rho_{\Lambda}$ is large, only small, non-zero values generate conditions in which many observers can exist. The condition set by Weinberg in Ref. [40] is that “observers require galaxies,” strengthened later to “observers require galaxies at least as large as the Milky Way,” which then frames the problem as using initial conditions and energy evolution to determine the density of large galaxies which potentially host the chemistry required for life. Bousso and collaborators find that an alternative approach is to measure significant entropy increase after decoupling, utilizing the idea that free energy is required for observation and so observation will scale with entropy production, following the second law of thermodynamics [39]. This technique proves successful in reproducing the observed $\rho_{\Lambda}$, and determines that most entropy production comes from IR photons produced by dust reprocessing of starlight, a process which corresponds to the presence of galaxies, stars, and heavy elements (all of which are required for life). Our results reaffirm that the largest source of post-decoupling entropy production from matter is from starlight reprocessing by dust.
@@
$s_{{\rm C}\nu{\rm B}}=\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{c^{3}\hbar^{3}}g_{\nu}T_{{\rm C}\nu{\rm B}}^{3}={\color[rgb]{0,0,0}(1.408\pm 0.002)}\times 10^{9}~k_{B}~\mathrm{m^{-3}}.$ (3.6)
with $g_{\nu}=6$ in the equation above for the 6 internal degrees of freedom of the three Standard Model neutrinos, and the (7/8) coefficient from the standard Fermi-Dirac correction. The error corresponding to the cosmic neutrino temperature is, as for the CMB, completely negligible (one in a billion). The corresponding entropy in the universe reads
@@
$\langle v\rangle\simeq 160\ (1+z)\ \frac{\rm km}{\rm sec}\ \left(\frac{\rm eV}{m_{\nu}c^{2}}\right).$ (3.8)
@@
$\lambda_{\rm fs}\simeq 2\ {\rm Mpc}\left(\frac{\rm eV}{m_{\nu}c^{2}}\right).$ (3.10)
@@ 4.1 Diffuse unbound baryons
The WHIM is the primary reservoir of baryonic matter ($\sim$40-50% of baryons, strongly supported by both simulations [58, 59, 60] and observations [61, 62, 63] and resides in filamentary structures of the cosmic web, with temperatures of 105-107 K [58, 59, 60, 64, 65]. UV/X-ray absorption studies (e.g., OVI, OVII, OVIII) confirm WHIM’s presence [61, 62, 66]. Stacking techniques ([63]) and thermal Sunyaev-Zeldovich (SZ) surveys ([67]) reveal WHIM in dense filaments, though detection of hotter phases ($10^{6}$ K) remains incomplete.
[…]
@@
ISM (Cold ISM) $10^{1}-10^{2}$ (1010010-100) $\sim 65$ $(5.7-57)\times 10^{80}$
ISM (Warm ISM) $\sim 10^{4}$ (0.110.1-1) 75-76 $(6.6-67)\times 10^{80}$
@@
We estimate the contribution to the total entropy of the universe of diffuse baryons using the ranges of density and temperature detailed upon above; Table 1 lists the entropy per baryon and the total entropy for each component, indicating that all components have comparable entropy contributions, with the IGM components contributing most likely the largest part to the entropy in diffuse backgrounds, roughly a factor 3 more than the ISM and 30 or more than the ICM. Overall, we find that, at most, at 95% C.L. given all uncertainties, diffuse baryons contribute to the entropy of the universe an amount
@@
We discuss below in Section 6.1 clustering enhancement for dark matter, where substructure exists down to a cutoff produced by the microscopic properties of the dark matter candidate, and where the dark matter is slated, or assumed, to be in virial equilibrium (with the temperature, therefore, fixed by the virial theorem applied to the given (sub-) structure mass. In the case of baryonic matter, we can directly measure, as discussed above, the temperature of the various, different components, effectively, therefore, accounting for clustering.
@@ 4.2 Baryons in stellar and compact objects
The population of stars and stellar remnants in the observable universe can be estimated by extrapolating detailed studies of the Milky Way, combined with broader astrophysical models and scaling relationships. The Milky Way, with approximately $10^{11}\ {\rm to}\ 4\times 10^{11}$ stars and a cumulative stellar mass of $\sim 5-6\times 10^{10}M_{\odot}$, serves as a representative template for large galaxies [75, 76, 77]. There are an estimated 2 trillion galaxies in the observable universe, spanning a range of morphologies, masses, and star formation histories, with a total stellar mass of approximately $10^{23}M_{\odot}$ spread across these systems [78]. By assuming the Milky Way’s stellar population as an approximate baseline for typical galaxies, we can scale its stellar and remnant populations to estimate the census of stars, white dwarfs, and neutron stars across the observable universe. This is the main goal of the present section. We will then utilize the entropy-per-baryon estimates of Ref. [16] to infer the global entropy contribution of all relevant stellar and stellar remnants in the universe.
@@
The stellar remnants formed from these stars are predominantly white dwarfs, which are the endpoints of stars with initial masses less than $\sim 8$ - $10M_{\odot}$ [81, 78]. The Milky Way is estimated to contain approximately $\sim 10^{9}-10^{10}$ white dwarfs, distributed across a range of cooling stages. Recent cooling models, accounting for core processes such as latent heat from crystallization and energy release during $^{22}\mathrm{Ne}$ sedimentation, suggest that older, “cold” white dwarfs ($\sim 1\times 10^{5}\,\mathrm{K}$) dominate the population, likely comprising $90\%$ of all white dwarfs [78, 82, 83]. This implies a cold white dwarf population on the order of $2-3\times 10^{21}$ in the universe, while younger, hot white dwarfs ($\sim 5\times 10^{8}\,\mathrm{K}$) constitute a smaller fraction (1-3%), totaling $\sim 10^{20}$ objects [78, 82, 83].
Neutron stars, formed via supernova explosions of stars in the $8-25M_{\odot}$ range, are significantly less common than white dwarfs due to the IMF bias toward lower-mass progenitors [75, 79]. Population synthesis models predict approximately $\sim 10^{8}$ neutron stars in the Milky Way, primarily cooling remnants at temperatures of $\sim 10^{7}\,\mathrm{K}$ [84, 85]. Scaling to the universe, the cold neutron star population is on the order of $\sim 2-3\times 10^{20}$. Conversely, newly formed, “hot” neutron stars $(\sim 10^{9}\,\mathrm{K})$ exist only briefly (for seconds to minutes post-supernova) and number $10^{3}$ in the Milky Way at any given time, extrapolating to $\sim 10^{15}$ in the observable universe [84, 86].
@@
[…]